by selling a towel for rs. 126.90, a man loses 6%, for how much should he sell to gain 4%

Answer by mananth(16946)

He loses 6% on the sale
..
SP = selling price
CP = cost price
..
SP = 94 when CP = 100
SP== 126.90 CP=?
..
CP= 126.90*100/94
=135.
..
He wants 4 % profit
..
4% of 135 = (4/100)*135= Rs 5.40.
Add it to CP
135+5.40 = 139.40. will be SP
...
m.ananth@hotmail.ca